40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 48 of 327  Not logged in ELOG logo
ID Date Author Type Category Subject
  14098   Mon Jul 23 09:58:52 2018 SteveSummaryVACRGA scan at day 6

 

 

Attachment 1: pd81-560Hz-d6.png
pd81-560Hz-d6.png
  14097   Sun Jul 22 14:01:07 2018 poojaUpdateCamerasDeveloping neural networks on simulated video

Aim: To develop a neural network that resolves mirror motion from video.

Since error was high for the same input as in my previous elog http://nodus.ligo.caltech.edu:8080/40m/14089

I modified the network topology by tuning the number of nodes, layers and learning rate so that the model fitted the sum of 4 sine waves efficiently, saved weights of the final epoch and then in a different program, loaded saved weights & tested on simulated video that's produced by moving beam spot from the centre of image by sum of 4 sine waves whose frequencies and amplitudes change with time.

Input : Simulated video of beam spot motion in pitch by applying 4 sine  waves of frquencies 0.2, 0.4, 0.1, 0.3 Hz  and amplitude ratios to frame size to be 0.1, 0.04, 0.05, 0.08. This is divided into train (0.4), validation (0.1) and test (0.5).

Model topology:

                                          Input               -->                  Hidden layer               -->                    Output layer                                  

                                                                                          8 nodes                                              1 node

Activation function:                                  selu                                             linear

Batch size = 32, Number of epochs = 128, loss function = mean squared error

Optimizer: Nadam ( learning rate = 0.00001, beta_1 = 0.8, beta_2 = 0.85)

Normalized the target sine signal of NN by dividing by its maximum value.

Plot of predicted output by neural network, applied input signal & residual error given in 1st attachment. These weights of the model in the final epoch have been saved to h5 file and then loaded & tested with simulated data of 4 sine waves with amplitudes and frequencies changing with time from their initial values by random uniform noise ranging from 0 to 0.05. Plot of predicted output by neural network, target signal of sine waves & residual error given in 2nd attachment. The actual signal can be got from predicted output of NN by multiplication with normalization constant used before. However, even though network fits training  & validation sets efficiently, it gives a comparatively large error on test data of varying amplitude & frequency.

Gautam suggested to try training on this noisy data of varying amplitudes and frequencies. The results using the same model of NN is given in Attachment 3. It was found that tuning the number of nodes, layers or learning rate didn't improve fitting much in this case.

 

 

Attachment 1: nn_simulation_2_normalized_mult_sin_nodes8_128epochs_lr0p00001_beta1_0p8_beta2_0p85_0p4train_0p1valid_marked.pdf
nn_simulation_2_normalized_mult_sin_nodes8_128epochs_lr0p00001_beta1_0p8_beta2_0p85_0p4train_0p1valid_marked.pdf
Attachment 2: nn_simulation_normalizedtarget_128epochs_mult_sin_load_wt_varyingtest_nodes8_lr0p00001_beta1_0p8_beta2_0p85_0p4train_0p1valid_marked.pdf
nn_simulation_normalizedtarget_128epochs_mult_sin_load_wt_varyingtest_nodes8_lr0p00001_beta1_0p8_beta2_0p85_0p4train_0p1valid_marked.pdf
Attachment 3: nn_simulation_2_normalized_varying_mult_sin_nodes8_128epochs_lr0p00001_beta1_0p8_beta2_0p85_0p4train_0p1valid_marked.pdf
nn_simulation_2_normalized_varying_mult_sin_nodes8_128epochs_lr0p00001_beta1_0p8_beta2_0p85_0p4train_0p1valid_marked.pdf
  14096   Sat Jul 21 14:03:19 2018 KojiSummaryThermal CompensationY arm locking

Ah. With MC2 feedback, we have about 3 times smaller "optical gain" for the ASS A2L. We have same dither, same actuator, but we need only 1/3 actuation of the MC2 compared to the ETMY case.
We had to reduce the ASS spot servo from 1 to 0.3 to make is stable, so this means that the ASS is really merginally stable.

  14095   Sat Jul 21 01:14:02 2018 gautamUpdateOMCPZT Jena driver board check

[Aaron, gautam]

We did a quick check of this board today. Main takeaways:

  • There are two voltages (HV pos and HV neg) that are output from this unit.
  • Presumably, these goto different piezoelectric elements, referenced to ground. Are there any spec sheets for these describing the geometry/threshold voltages?
  • The outputs are:
    • \mathrm{HV_{+}} = 10(V_{\mathrm{DAC}}+V_{\mathrm{offset}}), \mathrm{HV_{-}} = 10(-V_{\mathrm{DAC}}+V_{\mathrm{offset}})
    • So with V_{\mathrm{offset}} = 7.5 \mathrm{V}, we expect to be able to use +/- 7.5 V of DAC range.
  • The trim pot had to be adjusted to realize V_{\mathrm{offset}} = 7.5 \mathrm{V}​.
  • I assume 150V is some kind of damage threshold of the PZT, so there is no benefit to using 10V offset voltage (as this would result in 200 V at full range DAC voltages).

With the correct V_{\mathrm{offset}} = 7.5 \mathrm{V}, we expect 0V from the DAC to result in 0 actuation on the mirror, assuming that an equal 75V goes to 2 PZTs mounted diametrically opposite on the optic. Hopefully, this means we have sufficient range to scan the input pointing into the OMC and get some sort of signal in the REFL signal (while length PZT is being scanned) which indicates a resonance. 

We plan to carve out some IFO time for this work next week.

  14094   Sat Jul 21 01:06:49 2018 gautamSummaryThermal CompensationY arm locking

I implemented this today. For now, the LSC output matrix is set to actuate on MC2 for Y arm locking. As expected, the transmission was much more stable, and the PLL control signal RMS was also reduced by factor of ~3. MC2 control signal does pick up a large (~2000 cts) DC component over a few hours, so we need to relieve this periodically.

Now that we have a workable ASS for the Y arm as well, we should be able to have more confidence in returning to the same beam spot position on the ETM and staying there during a scan using this technique.

The main improvement to be trialled next in the scanning is to improve the speed of scanning. As things stand, my script takes ~2.5 seconds per datapoint. If we can cut this in half, that'd be huge. On Wednesday night, we were extraordinarily lucky to avoid MC3 glitching, EPICS/slow machine failures, and GPIB freezes. Today, the latter reared its head. Fortunately, since I'm dumping data to file for each datapoint, this means we at least have data till the GPIB freeze.

Quote:

For future measurements, we should consider locking the IMC length to the arm cavity - this would eliminate such alignment drifts, and maybe also make the PLL control signal RMS smaller. 


Not related to this work: Terra, Sandrine, Keerthana and I cleaned up the lab a bit today, and made better cable labels. Aaron and I have to clean up the OMC area a bit. Huge thanks to Steve for taking care of our mess elsewhere in the lab!

  14093   Fri Jul 20 22:53:15 2018 KojiUpdateASSAttempt to resurrect Yarm ASS

[Koji Gautam]

We managed to realize stable ASS configuration for Yarm. The transmission of 1.06~1.07 was recovered by introducing intentional beam spot offset in the horizontal direction towards the opposite side of the elliptic reflector. The end table optics were adjusted to have the spots about the center of the mirrors, lenses, and PDs/QPDs.


Preparation

- The Y arm was manually aligned with a given input axis. The transmission was ~0.8.
- Then, TT2 was moved in yaw such that it introduced the horizontal beam shift at the end. By moving the spot to the opposite side of the reflector. The transmission ~0.95 was obtained after patient alignment work.

- Went to the end table and checked the spots. The beam was not at the center of the last 1" lens for the Trans PDs. The beam steering was adjusted to have the spot nicely going through the lens and the mirrors. This made the transmission level to be ~1.05.

- The beam centering on the Trans PD was checked and adjusted.
- The beam centering on the RF BBPD for the arm scan was checked. The spot was too big for that PD. The lens was slightly moved away from the PD to make the spot on the BBPD small. Now the PD saw the plateu when the steering was scanned (i.e. the spot is small enough).

- With the Y arm locked with MC2, the servo gain needs to be 0.012 instead of nominal 0.015 with ETMY to prevent from servo oscilating.

ASS tuning

- First of all, only the bottom 4 loops out of total 8 loops were tuned. They are the servos for the beam alignment with regard to the caivty. The linearity and the zero crossings were checked with regard to the reference alignment. All of these 4 showed offsets that causes the servo running away. Don't know the reason of this offset, but it is freq dependent. Therefore the dither freqs were tuned to make the offset zeroed, and tuned the demod phases there. This kept the transmission as high as the reference (~1.05)

- This allowed us to play with the spot position a bit by tuning the caivty alignment. In the end, the transmission of ~1.08 was obtained. Using this alignment, A2L offset for ETMY Yaw was determined to be +17 (to make the error signal -17). This offset produces almost a beam radius (5mm) shifted on the end mirror towards the opposite direction of the reflector.

- The nominal servo setting made the spot servo running away. Gautam pointed out that this could be a gain hierarchy problem (i.e. the spot servos are too fast). We ended up reducing the gain of the servo from 1.0 to 0.3 to make the spot servo stable.

- All the ASS setting was stored in a new snap file "script/ASS/ASS-DITEHR_ON.snap". The previous snap was saved to "script/ASS/ASS_DITHER_ON_preVent201807.snap". This did not save the exc gains of the oscillators. Therefore "DITHER_ASS_ON.py" was modified to have the new exc gains (CLKGAIN). The old values are stored in the comments in this script.


Overall this is not an ideal situation as we don't know what is the actually cause of the offsets in the dither error signals. We expect to correct the beam clipping and the suspension sooner or later. Therefore, we will come back to the ASS again once the other issues are corrected.

Attachment 1: 02.png
02.png
  14092   Fri Jul 20 22:51:28 2018 KojiUpdateIOOIMC WFS path alignment

IMC WFS tuning

- IMC was aligned manually to have maximum output and also spot at the center of the end QPD.
- The IMC WFS spots were aligned to be the center of the WFS QPDs.
- With the good alignment, WFS RF offset and MC2 QPD offsets were tuned via the scripts.

  14091   Fri Jul 20 18:30:47 2018 JonConfigurationAUXRecommend to install AUX PZT driver

I recently realized that the PLL is only using about 20% of the available actuation range of the AUX PZT. The +/-10 V control signal from the LB1005 is being directly inputted into the fast AUX PZT channel, which has an input range of +/-50 V.

I recommend to install a PZT driver (amplifier) between the controller and laser to use the full available actuator range. For cavity scans, this will increase the available sweep range from +/-50 MHz to +/-250MHz. This has a unique advantage even if slow temperature feedback is also implemented. To sample faster than the timescale of most of the angular noise,  scans generally need to be made with a total sweep time <1 sec. This is faster than the PLL offset can be offloaded via the slow temperature control, so the only way to scan more than 100 MHz in one measurement is with a larger dynamic range.

  14090   Fri Jul 20 07:43:54 2018 SteveSummarySUSETMY

 

 

Attachment 1: ETMY_leveling.png
ETMY_leveling.png
Attachment 2: ETMY.png
ETMY.png
  14089   Thu Jul 19 18:09:17 2018 poojaUpdateCamerasUpdate in developing neural networks

Aim: To develop a neural network that resolves mirror motion from video.

Case 1:

Input : Simulated video of beam spot motion in pitch by applying 4 sine  waves of frquencies 0.2, 0.4, 0.1, 0.3 Hz  and amplitude ratios to frame size to be 0.1, 0.04, 0.05, 0.08

The data has been split into train, validation and test datasets and I tried training on neural network with the same model topology & parameters as in my previous elog (https://nodus.ligo.caltech.edu:8081/40m/14070)

The output of NN and residual error have been shown in Attachment 1. This NN model gives a large error for this. So I think we have to increase the number of nodes and learning rate so that we get a lower error value with a single sine wave simulated video ( but not overfitting) and then try training on linear combination of sine waves.

Case 2 :

Normalized the target sine signal of NN so that it ranges from -1 to 1 and then trained on the same neural network as in my previous elog with simulated video created using single sine wave. This gave comparatively lower error (shown in Attachment 2). But if we train using this network, we can get only the frequency of test mass motion but we can't resolve the amount by which test mass moves. So I'm unclear about whether we can use this.

Attachment 1: nn_simulation_mlt_sine_nodes4_lr0p00001_beta1_0p8_beta2_0p85_marked.pdf
nn_simulation_mlt_sine_nodes4_lr0p00001_beta1_0p8_beta2_0p85_marked.pdf
Attachment 2: nn_simulation_2_nodes4_target-1to1_marked.pdf
nn_simulation_2_nodes4_target-1to1_marked.pdf
  14088   Thu Jul 19 13:35:30 2018 SteveSummaryVACannuloses pumped

Roughing down the annuloses required closing V1 for 13 minutes

IFO is 2.2e-5 Torr

Attachment 1: AnsPumped.png
AnsPumped.png
  14087   Thu Jul 19 11:01:03 2018 SteveSummaryVACpd81 @ 2e-5 Torr

Cold cathode gauge just turned on.

Attachment 1: pd81@2days.png
pd81@2days.png
  14086   Thu Jul 19 04:44:09 2018 Annalisa, TerraSummaryThermal Compensationfrequency shift observed with heating!

Annalisa, Gautum, Koji, Terra

Summary: with the reflector setup, we measured a frequency shift of the first and second order modes! First looks of shifts show 1st HOM shift ~-10 kHz, 2nd HOM shift ~-18 kHz (carrier ~4 kHz). We saw no shift with the cylinder/lenses set up.

- - - - -

Tonight we modified the cavity scan setup: the LO is provided by the Marconi which, at the same time, is also used to scan the AUX laser frequency instead of the Agilent. In order to get rid of the free running noise between Marconi and Agilent, the Marconi frequency was scanned and, point by point, the Agilent center frequency was changed accordingly. In order to speed up the process, the whole procedure was automated. The script is called AGfast.py and can be found in /users/annalisa/postVent.

One thing that helped in improving the data quality of the phase information was to set the Agilent IF bandwidth @1kHz. Not yet clear why, but it was better than having a lower bandwidth. To be further investigated.

With this setup, we made some coarse scan of the full FSR and then we "zoomed" around the main peaks in order to increase the resolution and get a more precise information about the peak frequency.

Here are the frequency ranges that we scanned:

  • carrier - central frequency: 31.73MHz; range: [31.68MHz - 31.78MHz]
  • HOM1 - central frequency: 32.88MHz; range: [32.84MHz - 32.93MHz]
  • HOM2 - central frequency: 34.03MHz; range: [33.95MHz - 34.06MHz]
  • HOM3 - central frequency: 35.18MHz; range: [35.09MHz - 35.25MHz]

We powered the heater of the lenses setup @4:55 am at 14.4V and 0.9A. Then we slightly increased the power @5:05am and the final "hot state" configuration is with heater powered at 16V and 0.9A.

With this setup we couldn't see any frequency shift

Then, at around 6:30 am we turned on the reflector setup and we measured a frequency shift of the first and second order modes. First scans show 1st HOM shift ~10 kHz, 2nd HOM shift ~18 kHz. First attachment shows carrier hot/cold, second attachment shows HOM2 hot/cold. We started to get plauged by high seismic noise. Heaters turned off at 7:45 am. Lots of scans and actual analysis to go.


gautam: about the questionable plotting -

  • 10 faint (alpha~0.3) lines are individual measurements with the reflector doing its heating. (AG4395A, 0 span, single frequency measurements plotted together).
  • charcoal line, labelled mean, is the mean of the 10 above lines.
  • bright green ("Reference") is the mean of a coarse scan (cold ETM) overlaid for comparison. 
  • "cold" - self explanatory.

My personal favourite plot is Attachment #3, which is a 5 MHz scan (cold) to identify positions of the various peaks. The power of including phase information in the analysis is clear. The second FSR on the right edge of the plot is not as prominent as the first is because the arm transmission was degrading throughout the measurement. For future measurements, we should consider locking the IMC length to the arm cavity - this would eliminate such alignment drifts, and maybe also make the PLL control signal RMS smaller. 

Attachment 1: scanning_fine_2018-07-19-07-32-08_parsed.pdf
scanning_fine_2018-07-19-07-32-08_parsed.pdf
Attachment 2: scanning_fine_2018-07-19-06-57-47_parsed.pdf
scanning_fine_2018-07-19-06-57-47_parsed.pdf
Attachment 3: Yscan_scanning_parsed_2am.txt.pdf
Yscan_scanning_parsed_2am.txt.pdf
  14085   Thu Jul 19 01:56:25 2018 gautamSummaryVACAUX pump shutdown

[koji, gautam]

Per Steve's instructions, we did the following:

  • TP3fl pressure reading was 65 torr.
  • TP3 controller reported pumping current of ~0.18A, temperature of 24C.
  • We throttled the manual valve which was connecting the "AUX" pump to the TP3fl.
  • The TP3fl pressure went up to 330 torr.
  • TP3fl controller reported current of 0.22A, temperature of 24C.
  • After ~5mins, we shut the AUX pump off.
  • We have monitored it over the last 1hour, no red flags.
    • (Before stopping AUX RP)
      0:56AM TP3 I=0.18A, P=6W, 23degC, TP3FL: 66
    • 0:59AM TP3 I=0.22A, P=7W, 23degC, TP3FL: 336
    • 1:15AM TP3 I=0.21A, P=7W, 23degC, TP3FL: 320
    • 1:31AM TP3 I=0.21A, P=7W, 23degC, TP3FL: 310
    • 2:06AM TP3 I=0.21A, P=7W, 23degC, TP3FL: 301
    • 5:06AM TP3 I=0.21A, P=7W, 23degC, TP3FL: 275
  14084   Wed Jul 18 23:43:50 2018 KojiUpdateGeneralVent 80 recovery

Is the reflector too close to the beam and causing clipping?

Quote:

For unknown reasons, the Y arm ASS does not maximize TRY. So we are in the unfortunate situation of neither arm having a working ASS servo. To be worked on later.

Attachment 1: IMG_5868.JPG
IMG_5868.JPG
Attachment 2: IMG_5382.JPG
IMG_5382.JPG
  14083   Wed Jul 18 17:36:50 2018 SteveSummaryVACpumpdown 81 at 6 +9 hrs completed

IFO P1 6e-4 Torr,  manual gate valve is fully open

The annuloses will be pumped down tomorrow.

Valve configuration: vacuum normal, RGA and annuloses are not pumped

Quote:

The manual gate valve scan was clean. Atm1     TP1 was pumping on it overnight.

                                                Pumpdown continued to hand over the pumping to TP1 maglev turbo

V1 was opened at P1 400 mTorr  with manual gate at 3/4 turn open position as Magev at 560 Hz rotation.

This is the first time we pumping down from atm with one small "beer can" turbo  and throttled gate to control load on small turbo forepump

The 70 l/s turbo is operating at 50k RPM, 0.7 A and 31 C,  pumping speed  ~ 44 mTorr/h at 200-400 mTorr range.

Watching foreline pressures and current one can keep opening gate valve little by little the so the load is optimized. It is working but not fast.

Let's keep small turbo at 0.8 Amp and 32 C max at this pumpdown. 

Quote:

10:20PM

  • Opened VM2 to pump down the RGA section with TP1
  • Stopped rotary roughing pumps
    • Manually closed RV1
    • Closed V3
    • Stopped RP1 and RP3
    • Vented the RP hose

The P1 pressure is 380mTorr. I allowed Gautam to use the full PSL power (~1W).

Attachment 1: pd81completed.png
pd81completed.png
Attachment 2: pd81@30hrs.png
pd81@30hrs.png
  14082   Wed Jul 18 12:49:08 2018 SteveSummaryVACpumpdown 81 at 6 +4.5hrs

The manual gate valve scan was clean. Atm1     TP1 was pumping on it overnight.

                                                Pumpdown continued to hand over the pumping to TP1 maglev turbo

V1 was opened at P1 400 mTorr  with manual gate at 3/4 turn open position as Magev at 560 Hz rotation.

Two aux fans on to hold tubo temps TP1 & TP3 . Atm3

This is the first time we pumping down from atm with ONE small "beer can" turbo  and throttled gate valve to control load on small turbo forepump

The 70 l/s turbo is operating at 50k RPM, 0.7 A and 31 C,  pumping speed  ~ 44 mTorr/h at 200-400 mTorr range with aux drypump in the foreline of TP3

Watching foreline pressures and current one can keep opening gate valve little by little the so the load is optimized. It is working but not fast.

Let's keep small turbo at 0.8 Amp and 32 C max at this pumpdown. 

Quote:

10:20PM

  • Opened VM2 to pump down the RGA section with TP1
  • Stopped rotary roughing pumps
    • Manually closed RV1
    • Closed V3
    • Stopped RP1 and RP3
    • Vented the RP hose

The P1 pressure is 380mTorr. I allowed Gautam to use the full PSL power (~1W).

Attachment 1: manlGateScan.png
manlGateScan.png
Attachment 2: handing_over_Mag.png
handing_over_Mag.png
Attachment 3: TGVw2auxfans_.jpg
TGVw2auxfans_.jpg
  14081   Wed Jul 18 03:14:48 2018 AnnalisaUpdateGeneralVent 80 recovery

[Gautam, Johannes, Koji, Annalisa]

Tonight we increased the power of the PSL laser and we achieved the lock of both arms with high power.

The AUX beam alignment to the Y arm was recovered and the PLL restored (using the Marconi as LO).

We made a quick measurement of the phase noise and the results will be posted tomorrow.

The beam on the PSL has been blocked, as well as the AUX beam on the AS table. The Marconi has been switched off.


gautam:

  1. Before turning up PSL power, I placed a block in front of MC refl to avoid any PD burning. Replaced HR Y1 2" optic with the usual 10% reflective BS to direct MC REFL to the locking PD.
  2. Waveplate was rotated back to 180 deg (original position before the vent). After optimizing PMC transmission, I measured 1.05 W going into the IMC (pre-vent value was 1.07 W, prolly within power meter absolute accuracy).
  3. IMC autolocker restored to usual high power version on megatron.
  4. There seems to be some kind of vacuum interlock in effect that prevents me from opening the PSL shutter via EPICS - I had to toggle the position on the shutter controller under the table. After tonight's work, I returned the controller to the NC state, to avoid any further interference with this interlock code that may prevent pumping in the AM.
  5. PLL gain was re-adjusted to achieve maximum stability (judged by eye) of the beat-note in lock triggered on the Marconi LO signal. Alignment onto the NF beatPD was also tweaked to squeeze out as much beat as possible.
  6. The main objective tonight was to send AUX beam in, recover transmission beat, scan the AUX frequency, and resolve some peaks (MAX HOLD scanning technique, magnitude only for now, no phase info). Thanks to JE's expert fiber alignment and beatnote maximization, we achieved this yes. Annalisa will post a plot tmr. 
  7. For unknown reasons, the Y arm ASS does not maximize TRY. So we are in the unfortunate situation of neither arm having a working ASS servo. To be worked on later.
  14080   Tue Jul 17 22:25:41 2018 KojiSummaryVACpumpdown 81 at 6 hrs

10:20PM

  • Opened VM2 to pump down the RGA section with TP1
  • Stopped rotary roughing pumps
    • Manually closed RV1
    • Closed V3
    • Stopped RP1 and RP3
    • Vented the RP hose

The P1 pressure is 380mTorr. I allowed Gautam to use the full PSL power (~1W).

  14079   Tue Jul 17 18:16:38 2018 SteveSummaryVACpumpdown 81 at 6 hrs

Precondition:  4 days at atm.   Atm5

HEPA tent used during the vent at ETMY  It reduced partical count 10 fold of 0.5 and 0.3 micron particals. Atm6

New items in vacuum:  Clean manual gate valve [Cetec made] from John Worden with 6" id....as it came from Hanford... [ Throttle able gate valve- TGV ] Atm3

                                 ( note: we have 3 more identical in the lab. The original intention was to use them for purging gates )

                                  Optiform Au plated reflector , Induceramics heating elements, similar as existing Cooner cables and related lenses, hardwear. see 14078

                                  OMC related item : none......... 14,110

 

The pumpdown is at 510 mTorr with RP1 & RP3 still pumping. Koji will shut it down the roughing later tonight. Tomorrow morning I will start the pumping by switching over to TP1 maglev.

Thanks for Koji and Gautam'  help of the installation of the manual gate valve. Atm4  This will allow us to control the load on our Varian foreline D70 turbo TP3

 

Attachment 1: pd81@6hrs.png
pd81@6hrs.png
Attachment 2: before_c.jpg
before_c.jpg
Attachment 3: tgv_c.jpg
tgv_c.jpg
Attachment 4: TGVinstalled.jpg
TGVinstalled.jpg
Attachment 5: 4_days_vent.png
4_days_vent.png
Attachment 6: tentHEPA.jpg
tentHEPA.jpg
  14078   Tue Jul 17 17:37:46 2018 Annalisa, TerraConfigurationThermal CompensationHeaters installation

Summary

We installed two heaters setup on the ETMY bench in order to try inducing some radius of curvature change and therefore HOMs frequency shift.

------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------

We installed two heaters setup.

Elliptic reflector setup (H1): heater put in the focus of the elliptical reflector: this will make a heat pattern as descirbed in the elogs #14043 and #14050. 

 

Lenses setup (H2): heater put in a cylndrical reflector (made up with aluminum foil) 1'' diameter, and 2 ZnSe lenses telescope, composed by a 1.5'' and a 1'' diameter respectively, both 3.5'' focal length. The telescope is designed in such a way to focus the heat map on the mirror HR surface. For this latter the schematic was supposed to be the following: 

This setup will project on the mirror a heat pattern like this:

which is very convenient if we want to see a different radius of curvature for different HOMs. However, the power that we are supposed to have absorbed by the mirror with this setup is very low (order of 40-ish mW with 18V, 1.2A) which is probably not enough to see an effect. Moreover, mostly for space reasons (post base too big), the distances were not fully kept, and we ended up with the following setup:

In this configuration we won't probably have a perfect focusing of the heat pattern on the mirror.

In vacuum connections

See Koji's elog #14077 for the final pin connection details. In summary, in vacuum the pins are:

13 to 8 --> cable bunch 0

7 to 2 --> cable bunch 2

25 to 20 --> cable bunch 1

19 to 14 --> cable bunch 3

where Elliptic reflector setup (H1) is connected to cables 0 and 1, and the lenses setup is connected to cables 2 and 3.

Installed setup

This is the installed setup as seen from above:

Attachment 5: IMG_5380.JPG
IMG_5380.JPG
  14077   Tue Jul 17 12:55:45 2018 KojiSummaryGeneralStarted pumping

[Steve, Koji, Gautam]

We started pumping down at ~12:15PM.

Vent finalization ~ YEND

  • The table leveling was way off. This was adjusted by the balancing weight. (Attachment 1~3)
  • The alignment of ETMY was not too much off. Just aligned it with the oplev spot on MEDM and this already made the green flashing.
  • The Green TEM00 was maximized with ITMY and ETMY. This made the PSL IR flashing.
  • The heater wires were checked. I found that one of the heater wires was touching the optical table via the cable shield. This is because the upper pins were shifted to the left side (Attachment 4&5). The pins were shifted and now all 4 cables are isolated form the table. I also checked the mutual resistance between the 4 terminals. They were measured to be isolated except two pairs that showed 4.4 Ohms and 4.0 Ohms (Attachment 6)
  • The tools were removed from the chamber. The Y arm was still flashing.
  • We closed the ETMY door.

Vent finalization ~ Vertex

  • Found the ITMX stuck. Gautam came in and showed us his black magic to release the optic...
  • This allowed us to align X arm. The green flash was found and the TEM00 flash was seen. This allowed us to see the PSL IR flash at the X end.
  • PRM Refl was aligned. SRM was aligned with the oplev.
  • The beam on the AS port was checked. The AS beam came out from the window.
  • Closed the OMC chamber.

Pumping

  • Started pumping with RP1 and RP3. (~12:15PM)

Attachment 1: IMG_5408.JPG
IMG_5408.JPG
Attachment 2: IMG_5400.JPG
IMG_5400.JPG
Attachment 3: IMG_5401.JPG
IMG_5401.JPG
Attachment 4: IMG_5402.JPG
IMG_5402.JPG
Attachment 5: IMG_5403.JPG
IMG_5403.JPG
Attachment 6: IMG_5404.JPG
IMG_5404.JPG
  14076   Tue Jul 17 12:46:28 2018 ranaUpdateGeneralsome notes from yesterday

For the EY, instead of balancing the table, I just moved the weight approximately so that the ETMY OSEMS were at half light, but didn't check the level since ETMY is the only optic.

 

Some notes on OMC/AS work (Aaron/Gautam can amend/correct):

- Beam is now well centered in OMC MMT. Hits input coupling mirror and cleanly exits the vacuum to the AS table.

- Didn't see much on OMC trans, but PDs are good based on flashlight test.

- just before closing, re-aligned beam in yaw so that it gets close to the east screw on the input coupler. Aaron and I think we maybe saw a flash there with the OMC length PZT being driven at full range by a triangle wave.

- with OMC Undulators (aka tip/tilt PZT mirrors) energized, the beam was low on PZT1 mirror. We pitched ITMY by ~150 micro-rad and that centered the beam on PZT1 mirror. ITMY-OL is probably not better than 100 urad as a DC reference?

- We checked the range of Undulator 1 and we were getting ~5 mrad of yaw of the beam for the full range, and perhaps half of that in pitch. Rob Ward emailed us from Oz to say that the range is robably 2.7 mrad, so that checks out.

Even if the ITMY has to be in the wrong position to get the beam to the OMC, we can still do the heater tests in one position and then do the OMC checkout stuff in the other position.

Gautam suspects that there is a possible hysterical behaviour in the Undulators which is related to the MC3 glitching and the slow machine hangups and also possibly the illuminati.

[aaron]

-We noticed a ghost beam that from MC REFL (MMT2) that should be dumped during the next vent--it travels parallel to the OMC's long axis and nearly hits one of the steering mirrors for OMC refl.

-We measured the level of the table and found it ~3 divisions off from level, with the south end tilted up

-Gautam rotated and slightly translated OM5 to realign the optic, as expected. No additional optics were added.

-Gautam and I tested the TT piezo driver. We found that 3.6V on the driver's input gave 75V (of 150V) at the output, at least for yaw on piezo 1. However, as Gautam mentioned, during testing it seemed that the other outputs may have different (nonzero) offset voltages, or some hysterisis.

  14075   Tue Jul 17 01:07:40 2018 gautamUpdateSUSETMY EQ stops

For the heater setup on EY table, I EQ-stopped ETMY. Only the face EQ stops (3 on HR face, 2 on AR face) were engaged. The EY Oplev HeNe was also shutdown during this procedure. 

  14074   Mon Jul 16 18:12:00 2018 KojiUpdateVACAdding a manual gate valve between TP1 and V1/VM2

[Steve Koji]

We are in the process of adding a manual gate valve between TP1 (Osaka Maglev) and the other gate valves (I suppose V1 and VM2).
The work is still on going and we will continue to work on this tomorrow. Because this section is isolated from the main volume, this work does not hold off the possible rough pumping tomorrow morning.

The motivation of this work is as follows:
- Since TP2 failed, the main vacuum volume has been pumped down by TP1 and TP3. However TP3 is not capable to handle the large pressure difference at the early stage of the turbo pumping. This cause TP3 to have excessive heating or even thermal shutdown.
- The remedy is to put a gate valve between TPs and the main vacuum to limit the amount of gas flowing into the TPs. This indeed slows down the pumping speed of turbo, but this is not the dominant part of the pumping time.

Actual work:
- Comfirmed TP1 is isolated.
- Unscrewed the flange of TP1.
- Remove TP1. This required to lift up TP1 with some shim as the nuts interferes with the TP1 body. (Attachment1, 2, 3)
- Now remove 10inch flange adapter. (Attachment4)
-
Attach 10"-8" adapter and 8" rotational sleeve. (Attachment5)

Attachment 1: P_20180716_155413.jpg
P_20180716_155413.jpg
Attachment 2: P_20180716_155645.jpg
P_20180716_155645.jpg
Attachment 3: P_20180716_155738.jpg
P_20180716_155738.jpg
Attachment 4: P_20180716_162307.jpg
P_20180716_162307.jpg
Attachment 5: P_20180716_172000.jpg
P_20180716_172000.jpg
  14073   Mon Jul 16 15:07:19 2018 KojiSummaryVACOven C vent

[Steve Koji]

- Attachment1: Removed the thermal cap. Checked the temperature of the oven. It was totally cold.

- Attachment2: Confirmed the RGA section was isolated. The pumps for the RGA was left running.

- Attachment3: Closed the main valve. The pumps for the main volume was left running.

- Attachment4: Started removing the rid. This did not change the gause readings as they were isolated from the venting main volume.

- Attachment5: Opened the rid. Took the components out on a UHV foil bag. The rid was replaced but loosely held by a few screws with the old gasket, just to protect the frange and the volume from rough dusts.

Attachment 1: P_20180716_141512.jpg
P_20180716_141512.jpg
Attachment 2: P_20180716_141601.jpg
P_20180716_141601.jpg
Attachment 3: P_20180716_141610.jpg
P_20180716_141610.jpg
Attachment 4: P_20180716_141827.jpg
P_20180716_141827.jpg
Attachment 5: P_20180716_143901.jpg
P_20180716_143901.jpg
  14072   Sat Jul 14 16:04:34 2018 aaronUpdateOMCChecking OMC Electronics

Next check is the DCPD/OMMT Satellite Box

I traced a cable from the OMC electrical feedthrough flanges to find the DCPD/OMMT Satellite Box (D060105). I couldn't find the DCC number or mention of the box anywhere except this old elog.

Gautam and I supplied the box with power and tested what we think is the bias for the PD, but don't read any bias... we tracked down the problem to a suspicious cable, labelled.

We confirmed that the board supplies the +5V bias that Rich told us we should supply to the PDs.

We tested the TFs for the board from the PD input pins to output pins with a 100kHz low pass (attached, sorry no phase plots). The TFs look flat as expected. The unfiltered outputs of the board appear bandpassed; we couldn't identify why this was from the circuit diagram but didn't worry too much about it, as we can plan to use the low passed outputs.

Attachment 1: Screenshot_2018-07-14_17.53.40.png
Screenshot_2018-07-14_17.53.40.png
Attachment 2: Screenshot_2018-07-14_17.57.17.png
Screenshot_2018-07-14_17.57.17.png
  14071   Fri Jul 13 23:39:46 2018 AnnalisaConfigurationThermal CompensationThermal compensation setup - power supply

[Annalisa, Rana]

In order to power the heater setup to be installed in the ETMY chamber, we took the Sorensen DSC33-33E power supply from the Xend rack which was supposed to power the heater for the seismometer setup.

We modified the J3 connector behind in such a way to allow a remote control (unsoldered pins 9 and 8). 

Now pins 9 and 12 need to be connected to a BNC cable running to the EPICS.


RXA update: the Sorensen's have the capability to be controlled by an external current source, voltage source, or resistive load. We have configured it so that 0-5V moves the output from 0-33 V. There is also the possibility to make it a current source and have the output current (rather than voltage) follow the control voltage. This might be useful since out heater resistance is changing with temperature.

Attachment 1: IMG_2012.jpg
IMG_2012.jpg
Attachment 2: IMG_2013.jpg
IMG_2013.jpg
Attachment 3: 20180713_213818.jpg
20180713_213818.jpg
  14070   Fri Jul 13 23:23:49 2018 poojaUpdateCamerasUpdate in developing neural networks

Aim: To develop a neural network that resolves mirror motion from video.

I tried to reduce the overfitting problem in previous neural network by reducing the number of nodes and layers and by varying the learning rate, beta factors (exponential decay rates of moving first and second moments) of Nadam optimizer assuming error of 5% is reasonable.

Input:

32 * 32 image frames (converted to 1d array & pixel values of 0 to 255 normalized) of simulated video by applying sine signal to move beam spot in pitch with frequency 0.2Hz and at 10 frames per second.

Total: 300 cycles ,           Train: 60 cycles,    Validation: 90 cycles,    Test: 150 cycles

Model topology:

                                          Input               -->                  Hidden layer               -->                    Output layer                                  

                                                                                          4 nodes                                              1 node

Activation function:                                  selu                                             linear

Batch size = 32, Number of epochs = 128, loss function = mean squared error

Optimizer: Nadam

Case 1:

Learning rate = 0.00001,    beta_1 = 0.8 (default value in Keras = 0.9),  beta_2 = 0.85 (default value in Keras = 0.999)

Plot of predicted output by neural network, applied input signal & residual error given in 1st attachment.

Case 2:

Changed number of nodes in hidden layer from 4 to 8. All other parameters same.

These plots show that when residual error increases basically the output of neural network has a smaller amplitude compared to the applied signal. This kind of training error is unclear to me.

When beta parameters of optimizer is changed farther from 1, error increases.

Attachment 1: nn_simulation_2_nodes4_lr0p00001_beta1_0p8_beta2_0p85.pdf
nn_simulation_2_nodes4_lr0p00001_beta1_0p8_beta2_0p85.pdf
Attachment 2: nn_simulation_2_nodes8_lr0p00001_beta1_0p8_beta2_0p85.pdf
nn_simulation_2_nodes8_lr0p00001_beta1_0p8_beta2_0p85.pdf
  14069   Fri Jul 13 20:36:33 2018 KojiSummaryGeneralIn vac/In air heater wiring

I went to the Y-end and took more photos of the cable stand. These revealed that in-vac pin #13 is connected to the shield of the cable (P.2). This in-vac pin #13 corresponds to  in-air pin #1. So in the end, we bunch the pins in the following order.

In Air In Vac
Pin #2-7 Pin #12-7
Pin #8-13 Pin #6-1
Pin #14-19 Pin #25-20
Pin #20-25 Pin #19-14

 

Attachment 1: heater_wiring.pdf
heater_wiring.pdf heater_wiring.pdf heater_wiring.pdf
  14068   Fri Jul 13 18:01:13 2018 gautamUpdateGeneralLow power MC

After getting the go ahead from Steve, I removed the physical beam block on the PSL table, sent the beam into the IFO, and re-aligned the MC to lock at low power. I've also revived my low power autolocker (running on megatron), seems to work okay though the gains may not be optimal, but it seems to do the job for now. Nominal transmission when well aligned at low power is ~1200cts. I briefly checked Y arm alignment with the green, seems okay, but didn't try locking the Y arm yet. All doors are still on, and I'm closing the PSL shutter again while Keerthana and Sandrine are working near the AS table.

  14067   Fri Jul 13 17:13:45 2018 KojiUpdateGeneralVent objectives and prep

Notice: I removed these 75Ohm video cables.

Attachment 1: P_20180713_020603.jpg
P_20180713_020603.jpg
  14066   Fri Jul 13 16:26:52 2018 SteveUpdateVACVent 80 is completing...

Steve and Aaron,

6 hrs vent is reaching equlibrium to room air. It took 3 and a half instrument grade air cilynders [ AI UZ300 as labelled ] at 10 psi pressure. Average vent speed ~ 2 Torr/min

Valve configuration: IFO at atm and RGA is pumped through VM2 by TP1 maglev.

 

Attachment 1: @atm.png
@atm.png
Attachment 2: vent80_7h.png
vent80_7h.png
Attachment 3: ventregN2&Air_c.jpg
ventregN2&Air_c.jpg
  14064   Fri Jul 13 10:54:55 2018 aaronUpdateVACVent 80

[aaron, steve]

Steve gave me a venting tutorial. I'll record this in probably a bit more detail than is strictly necessary, so I can keep track of some of the minor details for future reference.

Here is Steve's checklist:

  • Check that all jam nuts are tightened
  • all viewports are closed
  • op levs are off
  • take a picture of the MEDM screens
  • Check particle counts
  • Check that the cranes work & wiped
  • Check that HV is off

Gautam already did the pre-vent checks, and Steve took a screenshot of the IFO alignment, IMC alignment, master op lev screen, suspension condition, and shutter status to get a reference point. We later added the TT_CONTROL screen. Steve turned off all op levs.

We then went inside to do the mechanical checks

  • N2 cylinders in the 40m antechamber are all full enough (have ~700psi/day of nitrogen)
  • We manually record the particle count
    • this should be <10,000 on the 0.5um particles to be low enough to vent, otherwise we will contaminate the system
    • note: need to multiply the reading on the particle counter by 10 to get the true count
    • the temperature inside the PSL enclosure should be 23-24C +/- 3 degrees
    • We recorded the particle counts at ~830 and ~930, and the 0.5um count was up to ~3000
  • We put a beam stop in front of the laser at the PSL table
  • Checked that all HV supplies are either off or supplying something in air
    • we noticed four HV supplies on 1X1 that were on. Two were accounted for on the PSL table (FSS), and the other two were for C1IOO_ASC but ran along the upper cable rack. We got ahold of Gautam (sorry!) and he told us these go to the TT driver on OMC_SOUTH, where we verified the HV cables are disconnected. We took this to mean these HV supplies are not powering anything, and proceeded without turning these HV off.
    • There are HV supplies which were all either off or supplying something in-air at: 1Y4, 1Y2, OMC N rack, 1X9 (green steering HV)
  • Checked that the crane works--both move up and down
    • vertex crane switch is on the wall at the inner corner of the IFO
    • y arm crane switch is on the N wall at the Y end
    • turn off the cranes at the control strip after verifying they work
  • While walking around checking HV, we checked that the jam nuts and viewports are all closed
    • we replaced one viewport at the x arm that was open for a camera

After completing these checks, we grabbed a nitrogen cylinder and hooked it up to the VV1 filter. Steve gave me a rundown of how the vacuum system works. For my own memory, the oil pumps which provide the first level of roughing backstream below 500mtorr, so we typically turn on the turbo pumps (TP) below that level... just in case there is a calibrated leak to keep the pressure above 350mtorr at the oil pumps. TP2 has broken, so during this vent we'll install a manual valve so we can narrow the aperture that TP1 sees at V1 so we can hand off to the turbo at 500mtorr without overwhelming it. When the turbos have the pressure low enough, we open the mag lev pump. Close V1 if things screw up to protect the IFO. This 6" id manual gatevalve will allow us throttle the load on the small turbo while the maglev is taking over the pumping  The missmatch in pumping speed is 390/70 l/s [ maglev/varian D70 ]  We need to close down the conductive intake of the TP1 with manual gate valve so the 6x smaller turbo does not get overloaded...

We checked CC1, which read 7.2utorr.

Open the medm c0/ce/VacControl_BAK.adl to control the valves.

Steve tells me we are starting from vacuum normal state, but that some things are broken so it doesn't exactly match the state as described. In particular, VA6 is 'moving' because it has been disconnected and permanently closed to avoid pumping on the annulus. During this v ent, we will also keep pumping on the RGA since it is a short vent; steve logged the RGA yesterday.

We began the vent by following the vacuum normal to chamber open procedure.

  1. VM1 closed
  2. We didn't open VM3, because we want to keep the RGA on
  3. Closed V1
  4. Connect the N2 to the VV1 filter
    1. first puged the line with nitrogen
    2. We confirmed visually that V1 is closed
  5. We opened VM2 to pump on the RGA with the mag lev pump.
    1. This is a nonstandard step because we are keeping the RGA pumped down.
    2. The current on TP3 is ~0.19A, which is a normal, low load on the pump
  6. VV1 opened to begin the vent at ~10:30am
    1. use crescent wrench to open, torque wrench wheel to close
    2. Keep the pressure regulator below 10 psi for the vent. We started the vent with about 2psi, then increased to 8psi after confirming that the SUS sensors looked OK.
  7. We checked the pressure plot and ITMX/ETMX motion to make sure we weren't venting too quickly or kicking the optics
    1. Should look at eg C1:SUS-ITMX_SENSOR_LL, as well as C1:Vac-P1_pressure
  8. Once the pressure reaches 25torr, we switched over to dry air
    1. wipe off the outside dolly wheels with a wet rag, and exit through the x-arm door to get the air. Sweep off the area outside the door, and wipe off new air containers with the rag.
    2. Bring the cylinder inside, get the regulator ready/purged, and swap relatively quickly.
    3. We increased the vent speed to 10psi. 
    4. Steve says the vents typically take 4 of 300 cf cylinders from Airgas "Ultra Zero" AI  UZ300 that contains 0.1 PPM of THC

Everything looks good, so I'm monitoring the vent and swapping out cylinders.

At 12:08pm, the pressure was at 257 torr and I swapped out in a new cylinder.

Steve: Do not overpressurize the vacuum envelope! Stop around 720 Torr and let lab air do the rest. Our bellows are thin walled for seismic isolation.

Attachment 1: vent80wtiptilts.png
vent80wtiptilts.png
  14063   Fri Jul 13 02:52:11 2018 gautamUpdateGeneralVent objectives and prep

[KA, GV]

  1. Arm scan setup moved from West side of PSL table to North side of AP table
    • Marconi is now providing the LO for the loop. A ~5m 50ohm BNC cable needs to be laid out from the PLL area to the NW corner of the AP table where the Agilent is for the TF measurement scheme which allows phase discrimination.
    • We took this opportunity to characterize/improve the setup a bit, and also clean up the 10s of cables around the PSL table. There were some Video cables (75ohm instead of 50ohm no) that were part of the scanning setup which we excised.
    • Mode matching onto beat PD on PSL table was improved - beat signal Vpp increased by factor of 2. PLL gain was adjusted accordingly. 
    • AUX power @AP table ~37mW (before recombination BS), ~3.7mW onto SRM, ~200uW onto ITMY.
    • We got a beatnote in transmission of ~120uV (rms?) after optimizing alignment into Y arm cavity with the AUX frequency on an arm FSR.
    • From Sandrine's calculations, I expected ~20mV of beat signal. We leave it to Annalisa and Terra to square that circle. Probably the MM into the arm isn't stellar either, and we didn't check the polarization of the aux beam (that it is matched to the IFO's p-pol).
    • The entire AUX injection chain needs careful characterization before
  2. Vent prep
    • Both arms were locked to IR, TRY maximized using ASS, TRX maximized by hand.
    • GTRY and GTRX were also maximized.
    • All test mass/PRM/SRM oplevs were centered with this "good" alignment.
    • PSL power into the IMC was cut from 1.07 W (measured after G&H mirror) to 97 mW, by rotating the waveplate immediately after the PSL (original angle 180deg, new angle 216 degrees).
    • PMC was locked. I did not need to change any of the gain settings.
    • 2" R=10% BS in the IMC REFL path was replaced with a Y2, so there is no MCREFL till we turn the power back up.
    • IMC was locked. I touched up my IMC low power autolocker, but this probably needs a bit more touching up tomorrow. MC has remained locked at low power for ~25mins, but as I typed this, I jinxed it and it lost the lock. anyway we have good alignment references.
    • PSL shutter will remain closed.

@SV, we are ready to vent tomorrow. Aaron is supposed to show up ~830am to assist.

Attachment 1: preVentStatus.png
preVentStatus.png
  14062   Fri Jul 13 00:15:13 2018 Annalisa, TerraConfigurationAUXY arm cavity scan

[Annalisa, Terra, Koji, Gautam]

Summary: We find a configuration for arm scans which significantly reduces phase noise. We run several arm scans and we were able to resolve several HOM peaks; analysis to come.

-------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------

As first, we made a measurement with the already established setup and, as Jon already pointed out, we found lots of phase noise. We hypothesized that it could either come from the PLL or from the motion of the optics between the AUX injection point (AS port) and the Y arm. 

  • We first characterized the PLL loop phase noise by comparing the beat signal against the Agilent reference signal, and we found that the beat had lots of phase noise with respect to the reference. Decreasing the PLL gain, we got rid of the phase noise in the beat signal.
  • Next, for the optical path length induced phase noise, we took the transfer function between TransMon and REFL signal rather than TransMon and Agilent reference signal. This takes advatage of the fact that the TransMon and REFL both see optical path length phase noise, which therefore gets canceled out in the transfer function. 

In this configuration, we were able to do arm scans where the phase variation at each peak was pretty clear and well defined. We took several 10MHz scan, we also zoomed around some specific HOM peak, and we were able to resolve some frequency split. 

We add some pictures of the setup and of the scan.

The data are saved in users/OLD/annalisa/Yscans. More analysis and plots will follow tomorrow. 

Attachment 1: IMG_6492.JPG
IMG_6492.JPG
Attachment 2: IMG_6494.JPG
IMG_6494.JPG
  14061   Thu Jul 12 23:59:14 2018 gautamUpdateGeneralVent objectives and prep

Vent objectives:

  1. Install radiative heater setup in EY chamber.
  2. Re-direct 50% of AS beam to OMC
    • First, we need to check if the OMC trans PDs work. Else, we have to consider using some in-air PDs for the transmon.
    • Check that OMC length PZT works.
    • Check that OMC steering PZTs work (Koji and I have used this in 2016 to check AS beam clipping so they should work).

We only anticipate opening up the IOO chamber and the EY chamber.


Vent preparation: see here.

  14060   Thu Jul 12 21:16:25 2018 aaronUpdateOMCChecking OMC Electronics

In preparation for tomorrow's vent, I'm checking some of the OMC-related electronics we plan to use.

First up is the HV Piezo Driver (D060283).

(well, technically the first up was the Kepco HV power supply... but I quickly tested that its output works up to 300V on a multimeter. The power supply for OMC-L-PZT is all good!)

According to the DCC, the nominal HV supply for this board is 200V; the board itself is printed with "+400V MAX", and the label on the HV supply says it was run at 250V. For now I'm applying 200V. I'm also supplying +-15V from a Tektronix supply.

I used two DB25 breakout boards to look at the pins for the DC and AC voltage monitors (OMC_Vmon_+/-, pins 1/6, and OMC_Vmon_AC+/-, pins 2 and 7) on a scope. I hooked up a DS345 function generator to the piezo drive inputn (pins 1,6). According to the 2013 diagram from the DCC, there is just one drive input, and an alternative "dither in" BNC that can override the DAC drive signal. I leave the alternative dither floating and am just talking to the DAC pins.

Aspects of the system seem to work. For example, I can apply a sine wave at the input, and watch on the AC monitor FFT as I shift the frequency. However, anything I do at DC seems to be filtered out. The DC output is always 150V (as long as 200V comes from the supply). I also notice that the sign of the DC mon is negative (when the Vmon_+ pin is kept high on the scope), even though when I measure the voltage directly with a multimeter the voltage has the expected (+) polarity.

A few things to try:

  • The DC_Readout electronics scheme on the wiki has separate oscillator and control inputs. This diagram has lied to us in the past and is older, and the traces on top of the breadboard seem to only go to pins 1 and 6, but I'm going to first try to apply a voltage across pins 2 and 7 in case there actually is a separate control I'm ignoring.
    • Driving on these pins seems to do nothing

On further investigation this was the key clue. I had the wrong DCC document, this is an old version of this board, the actual board we are using is version A1 of D060283-x0 (one of the "other files")

Gautam and Koji returned at this point and we started going through the testpoints of the board, before quickly realizing that the DC voltage wasn't making it to the board. Turns out the cable was a "NULL" cable, so indeed the AC wasn't passing. We swapped out the cable, and tested the circuit with 30V from the HV supply to trim the voltage reference at U14. The minimum voltage we could get is 5V, due to the voltage divider to ground made by R39. We confirmed that the board, powered with 200V, can drive a sine wave and the DC and AC mons behave as expected.

  14059   Thu Jul 12 16:18:22 2018 SteveUpdateVACVent preparations

We are getting ready to vent.

 

Attachment 1: before_Vent.png
before_Vent.png
Attachment 2: before_Vent_cond.png
before_Vent_cond.png
  14058   Thu Jul 12 15:15:47 2018 SandrineUpdate Beat Note Measurements for Cavity Scans

(Gautam, Sandrine)

We calculated the expected power of the beat note for Annalisa's Y arm cavity scans. 

Beat Note Measurement

We began by calculating the transmitted power of the PSL and AUX. We assumed that the input power of the PSL was 25 mW and the input power of the AUX was 250 uW. We also assumed a loss of 25 ppm for the ITM and ETM. We used T1 = 0.0138 and T2 = 25 x 10-6. 

P_{t} = \frac{t _{1}^{2}t_{2}^{2}}{1+r_{1}^{2}r_{2}^{2}-2r_{1}r_{2}}

t = \sqrt{T}          

r = \sqrt{1-T-L} = {\sqrt{R}}

The transmitted power of the PSL is approximately 100 uW, and the transmitted power of the AUX is approximately 0.974 uW. 

P_{t}^{PSL} = 100 uW                          P_{t}^{AUX} = 0.974 uW

The beat note was calculated with the following:

P_{beat} = 2\sqrt{P_{PSL}P_{AUX}} = 20 uW

The  expected beat note should be approximately 20 uW. 

  14057   Thu Jul 12 14:06:39 2018 keerthanaUpdateelogFinesse and Analytical solution - Comparison

I tried to compare the cavity scan data we get from the Finesse simulation and that we expect from the Analytical solution. The diagram of the cavity I defined in Finesse is given below along with the values of different quantities I used. For the analytical solution I have used two different equations and they are listed below.

Analytical 1 - Blue Graph

\phi = \frac {2.L.\Omega_1}{c}

t_{cav} = \frac{t_e. t_f \exp^{-i\frac{\phi}{2}}}{1- r_f. r_e \exp^{-i\phi} }

T_{cav} = \left|{t_{cav}} \right|^2

 

Analytical 2 - Red Graph

F = \frac {4. r_f.r_e}{(1-r_f.r_e )^2}

\phi = \frac {2.L.\Omega_1}{c}

T_{cav} = \left|{t_{cav}} \right|^2 = \frac {(t_e.t_f)^2}{(1 - r_f . r_e)^2} \frac{1}{1+F(\sin\frac {\phi}{2})^2}

The graph obtained from both these solutions completely matches with each other.

Finesse Solution

The cavity which I defined in Finesse is shown below. The solution from Finesse and the Analytical solution also matches with each other. Another plot is made by taking the difference between Finesse solution and Analytical solution. The difference seems to be of the order of \approx 10^{-19}.

The Difference plot is also attached below.

Attachment 1: finesse_cavity.png
finesse_cavity.png
Attachment 2: Analytical1.pdf
Analytical1.pdf
Attachment 3: Finesse_Analytical.pdf
Finesse_Analytical.pdf
Attachment 4: Difference.pdf
Difference.pdf
  14056   Thu Jul 12 12:26:39 2018 aaronUpdateGeneralOMC revival

We found a diagram describing the DC Readout wiring scheme on the wiki page for DC readout (THIS DIAGRAM LIED TO US). The wiring scheme is in D060096 on the old DCC.

Following this scheme for the OMC PZT Driver, we measured the capacitance across pins 1 and 14 on the driver end of the cable nominally going to the PZT (so we measured the capacitance of the cable and PZT) at 0.5nF. Gautam thought this seemed a bit low, and indeed a back of the envelope calculation says that the cable capacitance is enough to explain this entire capacitance.

Gautam has gone in to open up the HV driver box and check that the pinout diagram was correct. We could identify the PZT from Gautam's photos from vent 79, but couldn't tell if the wires were connected, so this may be something to check during the vent.

UPDATE:

Turns out the output was pins 13 and 25, we measured the capacitance again and got 209nF, which makes a lot more sense.

  14055   Thu Jul 12 11:13:39 2018 gautamUpdateGeneralOMC revival

Aaron and I are going to do the checkout of the OMC electronics outside vacuum today. At some point, we will also want to run a c1omc model to integrate with rtcds. Barring objections, I will set up this model on one of the spare cores on the physical machine c1ioo tomorrow.

  Draft   Wed Jul 11 18:13:19 2018 keerthanaSummaryAUXGouy Phase Measurements from AUX-Laser Scans

From the Measurement Jon made, FSR is 3.967 MHz and the Gouy phase is 52 degrees. From this, the length of the Y-arm cavity seems to be 37.78 m and the radius of curvature of the mirror seems to be 60.85 m.

 

Guoy Phase = \cos^{-1} \sqrt{g1.g2}

\\ g = 1- \frac{L}{R}

L = \frac {c} {2*FSR}

FSR = Free spectral Range

L = Lenth of the arm

R = Radius of curvature of the mirror (R1 =\infty  , R2= unknown)

Quote:

This note reports analysis of cavity scans made by directly sweeping the AUX laser carrier frequency (no sidebands). The measurement is made by sweeping the RF offset of the AUX-PSL phase-locked loop and demodulating the cavity reflection/transmission signal at the offset frequency.

Y-Arm Scan

Due to the simplicity of its expected response, the Y-arm cavity was scanned first as a test of the AUX hardware and the sensitivity of the technique. Attachment 1 shows the measured cavity transmission with respect to RF drive signal.

The AUX laser launch setup is capable of injecting up to 9.3 mW into the AS port. This high-power measurement is shown by the black trace. The same measurement is repeated for a realistic SQZ injection power, 70 uW, indicated by the red curve. At low power, the technique still clearly resolves the FSR and six HOM resonances. From the identified mode resonance frequencies the following cavity parameters are directly extracted.

YARM Gautam V. Finesse Model Actual
FSR 3.966 MHz 3.967 MHz
Gouy phase 54.2 deg 52.0 deg

 

 

  14053   Wed Jul 11 16:50:34 2018 poojaUpdateCamerasUpdate in developing neural networks

Aim: To develop a neural network that resolves mirror motion from video.

I had created a python code to find the combination of hyperparameters that trains the neural network. The code (nn_hyperparam_opt.py) is present in the github repo. It's running in cluster since a few days. In the meanwhile I had just tried some combination of hyperparameters.

These give a low loss value of approximately 1e-5 but there is a large error bar for loss value since it fluctuates a lot even after 1500 epochs. This is unclear.

Input: 64*64 image frames of simulated video by applying beam motion sine wave of frequency 0.2Hz and at 10 frames per sec. This input data is given as an hdf5 file.

Train : 100 cycles,  Test: 300 cycles, Optimizer = Nadam (learning rate = 0.001)

Model topology:

                    256       ->      128    ->       1

Activation :        selu     selu           linear

Case 1: batch size = 48, epochs = 1000, loss function = mean squared error

Plots of output predicted by neural network (NN) & input signal has been shown in 1st graph & variation in loss value with epochs in 2nd graph.

Case 2: batch size = 32, epochs = 1500, loss function = mean squared logarithmic error

Plots of output predicted by neural network (NN) & input signal has been shown in 3rd graph & variation in loss value with epochs in 4th graph.

 

 

 

Attachment 1: graphs.pdf
graphs.pdf graphs.pdf graphs.pdf graphs.pdf
  14052   Wed Jul 11 16:23:21 2018 aaronUpdateOMCCoordination of the Output Mode-cleaner Mirror Insertion Expedition (COMMIE)

I started this document on my own with notes as I was tracing the beam path through the output optics, as well as some notes as I started digging through the elogs. Let's just put it here instead....

  1. Beam from AS port into OMMT
  2. Reflect off OM5-PJ
    1. TO DO: check that the PZT works
    2. 40/P/F/L, 1525-45-P
  3. Pick off from OMPO
    1. TO DO: determine how much power is needed for the pick off, choose an appropriate optic (for this vent probably 50-50 is fine)
    2. The PO beam goes to OM6
  4. Reflect off MMT1???
    1. TO DO: determine if this mirror has a PZT, get it working
      1. Has a PZT?
      2. Which PZT channel on the DAQ?
      3. Is there a cable going to from the DAC to the PZT?
      4. Is the PZT functional?
      5. How many PZTs does this mirror actually have?
    2. TO DO: determine the real name of this optic, find its recent history in the elog
    3. TO DO: determine the correct telescope parameters to optimally couple into the mode cleaner given the following:
    4. TO DO: look up how the radius of curvature (RC) of the OMC has changed, and therefore what telescope parameters are necessary
  5. Focused by MMT2???
    1. TO DO: determine if this mirror has a PZT
      1. Has a PZT?
      2. Which PZT channel on the DAQ?
      3. Is there a cable going to from the DAC to the PZT?
      4. Is the PZT functional?
      5. How many PZTs does this mirror actually have?
    2. TO DO: determine the real name of this optic, find its recent history in the elog
    3. TO DO: what about this optic is tunable? It looks bulky
  6. Columnated by MMT3???
    1. TO DO: determine if this mirror has a PZT
      1. Has a PZT?
      2. Which PZT channel on the DAQ?
      3. Is there a cable going to from the DAC to the PZT?
      4. Is the PZT functional?
      5. How many PZTs does this mirror actually have?
  7. Steered by MMT4???
    1. TO DO: determine the real name of this optic
    2. TO DO: why is this optic so small? Looks different from the rest, maybe weird space constraint
  8. Steered by MMT5???
    1. TO DO: why is this optic so large compared to OMMT4?
    2. TO DO: is there a more space efficient way of steering this beam, or even some way that avoids having to steer with three distinct optics
  9. Steered by MMT6???
    1. TO DO: Can this optic be removed with some clever new beam path?
  10. Cleaned by the OMC
    1. TO DO: Where does the promptly reflected beam from OMC1 go after it exits the chamber?
    2. TO DO: check the PZTs
      1. Has a PZT?
      2. Which PZT channel on the DAQ?
      3. Is there a cable going to from the DAC to the PZT?
      4. Is the PZT functional?
      5. How many PZTs does the OMC actually have?
    3. TO DO: Determine if a new OMC configuration would be more ideal for the squeezing experiment
      1. This is a large task, not part of this immediate vent
    4. TO DO: What is done with the OMC reflection? What is done with the transmission?
    5. TO DO: Check the logs about how the OMC had been in use; should be mostly from rob ward
  11. Reflected beam goes to the next chamber
  12. Transmitted beam is split by OM7???
    1. TO DO: find the actual name of this optic
    2. TO DO: why does this have the R/T that is does?
  13. Reflected beam goes to my OMPD
    1. TO DO: figure out what this PD is used for, and whether we even need it
      1. I think this might be the camera mentioned in 40m elog 21
      2. Elog 42 says the 4 QPDs for the OMC have meds screens located under C2TPT—is this a clue for channel names?
  14. Transmitted beam is reflected to the next chamber by OM8???
    1. TO DO: determine the name of this optic
    2. TO DO: Where does this beam go? What is it used for?
  15. Beam Dumps to add
    1. Transmission through OM5? Probably don’t need…
    2. OMMT1 transmission
    3. OMMT steering mirror transmissions
    4. OMC transmissions? Probably not?
    5. OMPD transmission?
    6. OM8 transmission
    7. Green scattering off of the window where the beam goes after GR_SM5
    8. Backscatter from the OMC prompt reflection to the window
    9. Backscatter from the OMC reflection to the window
    10. Backscatter from the MC beam off the window (this beam just travels through this chamber, interacts with no optics; there is also what looks like a small blue beam on this diagram, so maybe need to dump that backscatter too)
    11. Backscatter from the PO beam from OM6 going through the chamber window
    12. Backscatter from IM1 out the window
    13. There is a small blue beam from OMMT3 that goes through this window as well, I’m not sure exactly what is is from or for, or if it is physical (there are a few of these strange blue lines, i'm probably just misreading the diagram)
  16. TESTS TO DO
    1. Characterize the PZT control
    2. Lock the OMC with a PZT dither lock
      1. Eg elog 59
    3. “Tap-tap-tappy-tap test” to find resonanes
      1. Look at combination of PDH error signal and DCPD signal???
      2. See elog 86 for results from initial OMC install—Nov 2007
    4. Check wiggling wires, etc
    5. TFs to check? Vertical TF?
    6. OMC Length check— see for eg elog 768
  17. ADDITIONAL TO DO
    1. Mode matching calculation for new radius of curvature optics—see elog 1271
      1. The current MMT is not the optimal configuration even for the old Rc (see 3077 and 3088)

 

Notes during reading elog

  • Entry 590 has a labelled picture of the optics setup with OMC
  • Mention of omcepics at elog 894
  • Some important changes happened in elog 1823
    • 1''->2'' mirror out of the vacuum--I should check whether this is still there, or if it has been moved
    • [many more changes.....]
  • There were at one time 2 cameras monitoring OMCT and R (see 4492, 4493)
  • Some OMC PZT HV supply info is at elog 4738, 4740... 
  • There are some photos of the OMC table at elog 5120, and a note about moving some optics
  • Not strictly about the OMC, but I really like Suresh's diagram 6756, I'll make something similar for the OMC electronics
    • although it is about adding the tip tilt electronics, which I think required a new flange for the OMC chamber
  • OMC stage 1 and 2 are the steering mirrors going into the OMC, and were controlled by EPICS chans (6875, 6884)
    • these PZT HV supplies lived in OMC_SOUTH (or maybe 1Y3? see elog 6893), the driver in OMC_NORTH (LIGO-D060287)
    • Photos of these supplies in 7696
  • There are pictures of the OMC and its PZTs in 7401
  • The OMC HV supply was moved to power a different set of PZTs (see 7603)
  • Talk of replacing the PZTs with picomotors or tip/tilts in 7684
  • More pictures of the OMC table before the OMC was 'removed' are here (8114) and in 12563/12571 Gautam links to a Picassa album with pictures from just before the beam was diverted
  14051   Wed Jul 11 15:57:00 2018 aaronUpdateOMCReviving OMC electronics
Gautam showed me the electronics racks for the OMC PZTs and DAQ. I'm in the process of chasing down what channels we need, and confirming that we'll be able to plug the old antialiasing/imaging boards into the current DAC/ADC boards. I found what I think was Rob Ward's simlink model for the omc, located at
 
/cvs/cds/caltech/cds/rward-advLigo/src/epics/simLink/omc.mdl
 
Channels in this model:
  • 27 or 29 total ADC channels are used (depending whether we keep 2 spare adc chans)
    • 4 each go to ASC_QPD1/2 (8 chans total)
    • 5 go to TRANS_PD1, TRANS_PD2, REFL_PD, TRANS_PD1_UF, TRANS_PD2_UF. These PD are used for ASC and LSC.
    • 2 go to the LSC, one each for DVMDC, DVMAC, X3DC, and X4DC
    • 12 go to the ASC_PZT
    • 2 go to the SPARE_ADC (not sure what this is)
    • I think these channels are (or were at some point) defined in memory by /cvs/cds/caltech/chans/ipc/G1.ipc
      • I found this from elog 2860; it mentions that these should eventually be migrated over to a file C1.ipc, but I don't see any OMC channels in that file or any of the 'old' C1.ipc files, so I suppose it never happened or they were removed later
    • During this vent, we won't have ASC, so
  • 10 or 14 DAC channels are used (depending whether we keep 4 spare dac chans)
    • 2 from the LSC, one for CLK_OUT and one for "LSC"
    • 8 from ASC, including P1A, P1B, P2A, P2B, P1OSC, Y1OSC, P2OSC, Y2OSC
    • I think these channels are (or were at some point) saved to frames due to /cvs/cds/caltech/chans/daq/C1OMC.ini, which I found from elog 2073
    • At some point, the 33MHz mod depth was controlled by one of the spare OMC chans, C1:OMC-SPARE_DAC_CH_15. See elog 2126. I assume this is no longer the case, since c1omc is defunct.
    • Durnig this vent, we won't have ASC and don't need to CLK_OUT the LSC, so we may just need one DAC channel

As of at least Nov 2009, the .par file for the OMC was located at /cvs/cds/gds/param/tpchn_C2 (see elog 2316)

 
Electronics inventory:
  • Kepco HV supply, "OMC-L-PZT", labels indicate it goes to 250V, needs to be tested  ("TESTED OK 2014OCT12")
  • Tip/Tilt Piezo Driver, LIGO D060287
  • HV Piezo Driver, LIGO D060283
  • QPD Whitening Board, D060214
  • LIGO D050374/D050387
  • LIGO D050368/D050373

Need to check:

  • Can the ADC/DAC adapter boards (eg D0902006) drive whatever ~10V control signal we need across ~10m of SCSI cable?
  •  
  14050   Tue Jul 10 23:44:23 2018 AnnalisaConfigurationThermal CompensationHeater setup assembly

[Annalisa, Koji]

Today both the heater and the reflector were delivered, and we set down the setup to make some first test.

The schematic is the usual: the rod heater (30mm long, 3.8 mm diameter) is set inside the elliptical reflector, as close as possible to the first focus. In the second focus we put the power meter in order to measure the radiated power. The broadband power meter wavelength calibration has been set at 4µm: indeed, the heater emits all over the spectrum with the Black Body radiation distribution, and the broadband power meter measures all of them, but only starting from 4µm they will be actually absorbed my the mirror, that's why that calibration was chosen.

We measured the cold resistance of the heater, and it was about 3.5 Ohm. The heater was powered with the BK precision DC power supply 1735, and we took measurements at different input current.

Current [A] Voltage [V] Measured radiated power [mW] Resistance [Ohm]
0.5 2.2 20 4.4
0.8 6 120 7.5
1 11 400 11
1.2 18 970 15

We also aimed at measuring the heater temperature at each step, but the Fluke thermal camera is sensitive up to 300°C and also the FLIR seems to have a very limited temperature range (150°C?). We thought about using a thermocouple, but we tested its response and it seems definitely too slow. 

Some pictures of the setup are shown in figures 1 and 6.

Then we put an absorbing screen in the suspension mount to see the heat pattern, in such a way to get an idea of the heat spot position and size on the ETMY. (figure 2)

The projected pattern is shown in figures 3-4-5

The optimal position of the heater which minimizes the heat beam spot seems when the heater inserted by 2/3 in the reflector (1/3 out). However, this is just a qualitative evaluation.

Finally, two more pictures showing the DB connector on the flange and the in-vacuum cables.

Some more considerations about in-vacuum cabling to come.

Steve: how are you going to protect the magnets ?

Attachment 1: IMG_1992.jpg
IMG_1992.jpg
Attachment 2: IMG_2002.jpg
IMG_2002.jpg
Attachment 3: IR20180710_0364a.png
IR20180710_0364a.png
Attachment 4: IR20180710_0368.png
IR20180710_0368.png
Attachment 5: IR20180710_0360.png
IR20180710_0360.png
Attachment 6: IMG_1993.jpg
IMG_1993.jpg
Attachment 7: IMG_5322.JPG
IMG_5322.JPG
Attachment 8: IMG_5321.JPG
IMG_5321.JPG
  14049   Tue Jul 10 16:59:12 2018 Izabella PastranaHowToComputer Scripts / ProgramsTaking Remote TF Measurements with the Agilent 4395A

I copied the netgpibdata folder onto rossa (under the directory ~/Agilent/), which contains all the necessary scripts and templates you'll need to remotely set up, run, and download the results of measurements taken on the AG4395A network analyzer. The computer will communicate with the network analyzer through the GPIB device (plugged into the back of the Agilent, and whose communication protocol is found in the AG4395A.py file in the directory ~/Agilent/netgpibdata/).

The parameter template file you'll be concerned with is TFAG4395Atemplate.yml (again, under ~/Agilent/netgpibdata/), which you can edit to fit your measurement needs. (The parameters you can change are all helpfully commented, so it's pretty straightforward to use! Note: this template file should remain in the same directory as AGmeasure, which is the executable python script you'll be using). Then, to actually set up, run, and download your measurement, you'll want to navigate to the ~/Agilent/netgpibdata/ directory, where you can run on the command line the following: python AGmeasure TFAG4395Atemplate.yml

The above command will run the measurement defined in your template file and then save a .txt file of your measured data points to the directory specified in your parameters. If you set up the template file such that the data is also plotted and saved after the measurement, a .pdf of the plot will be saved along with your .txt file.

Now if you want to just download the data currently on the instrument display, you can run: python AGmeasure -i 192.168.113.105 -a 10 --getdata

Those are the big points, but you can also run python AGmeasure --help to learn about all the other functions of AGmeasure (alternatively, you can read through the actual python script).

Happy remote measuring! :)

 

 

 

 

  14048   Tue Jul 10 14:20:09 2018 steveUpdateGeneralprojector light bulb replaced

Bulb replaced at day 110  We have now spare now.

 

ELOG V3.1.3-